I shoot 79.4% from the free throw line. I am going to shoot 98 free throws. What is the standard deviation of my expected value?
Answer #1

X B ( n = 98 , P = 0.794 )


Mean = n * P = ( 98 * 0.794 ) = 77.812
Variance = n * P * Q = ( 98 * 0.794 * 0.206 ) = 16.0293


Standard deviation = (variance) = (16.0293) = 4.0037